OpenStudy (anonymous):

What are the differences between (x+y)^3 and x^3+y^3?

OpenStudy (anonymous):

(x+y)^3 you are multiplying (x+y) by itself 3 times and x^3+y^3 is multiplying each individual variable 3 times

OpenStudy (anonymous):

because first one is (x+y)(X+Y)(X+Y) and second is x^3+y^3

OpenStudy (anonymous):

Thank you for all your help it is greatly appreciated.

OpenStudy (anonymous):

(x+y)^3= x^3+y^3?+3xy(x+Y) here in the expression (x+y)^3 already x^3+y^3 is exist and extra 3xy(x+y) is added also........ second method.......... let x=2 and y=3 then (x+y)^3=(2+3)^3=5^3=125 and in x^3+y^3=2^3+3^3=8+27=35 here 125 is not equal to 35 so they are different

OpenStudy (anonymous):

thanks